Fermi National Laboratory


Accelerator Update

The integrated luminosity for last week was 7.4 pb-1.

Wednesday August 27
The day shift began with Operations monitoring store 2953, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

Tevatron (TeV) experts began a cogging study at 11:21 AM.

Operations terminated store 2953 at 12:20 PM.

Pbar resumed stacking at 2:05 PM.

Operations put the TeV through a dry squeeze at 2:19 PM.

Operations put the TeV into shot setup at 2:52 PM and established store 2956 at 4:41 PM. Its initial luminosity was 4.1E31. Pbar resumed stacking.

At 7:16 PM, the TeV suffered a sector B0 quench. Operators called in an expert and he discovered a bad VFC card for a quadrupole power supply (B0Q5).

The cryo system for sector A4 was recovering slowly. Operations contacted the Cryo technicians at CHL and they took over the cooldown. The techs found and replaced a bad actuator card.

Thursday August 28
The midnight shift began with the TeV turn on continuing, with Pbar stacking, and with MiniBooNE taking beam.

At 12:52 AM, the TeV suffered a sector B0 quench on the first ramp. Operations thought that maybe the cryo system hadn't stabilized.

A Pbar expert began a study period at 1:40 AM.

TeV expert began tuning the machine for studies at 3:20 AM.

Pbar studies finished at 6:38 AM.

At 7:00 AM, Operations began turning everything off.

Operations made Linac ready for access at 7:03 AM.

By 8 AM, nearly all areas were ready for access.

The Linac and Booster accessed ended and Operations was turning these machines back on by 11:53 AM. Operations had beam to Booster by 12:38 PM.

Booster experts began a study period at 1:10 PM.

By 6:47 PM, Operations began the Main Injector (MI) turn-on, but problems at 7:17 PM halted the this. An expert had to come in and replaced equipment.

Operations established beam to MiniBooNE at 8 PM.

Pbar began stacking at 9:38 PM, but the beam looked bad.

MI work finished and normal stacking resumed at 10:35 PM.

MiniBooNE was suffering from low beam rates at 11:11 PM. Operations called in experts to investigate.

Friday August 29
The midnight shift began with Pbar stacking, with the TeV turn-on continuing, and with MiniBooNE taking beam, but still having problems.

At 12:51 AM, a MininBooNE beam-monitoring device (TOR860) failed. Operations turned off their beam and contacted experts. Experts switched the device to a new channel at 1:19 AM.

Operations put the TeV through its first ramp at 2:44 AM.

Operations had trouble turning on the Tevatron Electron Lens at 3:39 AM. An expert came in and had it operating by 3:58 AM.

TeV expert began tuning the TeV for studies at 4 AM.

Today's Plan
The plan for today is to continue with TeV studies until around noon and then go into shot setup. The plan for the weekend is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/29/2003   email Fermilab